OpenStudy (anonymous):

Find the values of a and b that make f continuous everywhere. F(x)= (x^2-4)/(x-2) x< 2 ax^2-bx+2 2 (uncluding 2) < x < 3 4x-a+b x>3 (including 3)

OpenStudy (anonymous):

do you know \[\lim_{x\to 2}\frac{x^2-4}{x-2}\]?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

4

OpenStudy (agent0smith):

For the function to be continuous, all the endpoints must meet up. When x=2, the (limit of the) first function equals 4. So ax^2-bx+2 must equal 4. Then, when x=3, ax^2-bx+2 must equal 4x-a+b.

OpenStudy (agent0smith):

Plug all the x's in and you should be able to solve some simultaneous equations.

OpenStudy (anonymous):

so, for ax^2-bx+2.. I substitute 2 and 3 ..

OpenStudy (anonymous):

and then when i substitute 3, and get my result, the last equation must equal my result from the last one.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[4a+2b-2=4\] is one equation

OpenStudy (anonymous):

\[9a-3b+2=12-a+b\] is the other one

OpenStudy (anonymous):

but everything has variables, do they stay like that?

OpenStudy (agent0smith):

You have two equations and two variables, which you can solve simultaneously, just like if you had something like x+y = 2 3x+2y=6

OpenStudy (anonymous):

ok. so I have being sitting here, trying to do the math, but I still don't understand where I'm getting at. I mean the second equation, I got 4a-2b+2=4 , do I pick a variable and solve for it?

OpenStudy (agent0smith):

Use simultaneous equations... remember those from algebra?

OpenStudy (agent0smith):

You have two equations which can be used to solve for a and b.